OpenStudy (anonymous):

lim as x approaches positive infinity of (1 + 1/4x) ^ 3x

OpenStudy (anonymous):

make a guess, i bet it is right

OpenStudy (anonymous):

if you guess wrong we can work through it

OpenStudy (anonymous):

is it e or positive infinity??

OpenStudy (anonymous):

closer with the \(e\) but not exactly \(e\) but rather \(e\) so some power

OpenStudy (anonymous):

first off i assume it is \[\lim_{x\to \infty}\left(1+\frac{1}{4x}\right)^{3x}\] right?

OpenStudy (anonymous):

probably taught to "take the log" get \[3x\ln(1+\frac{1}{4x})\] then it is in the form \(\infty\times 0\) usual gimmick is to rewrite as \[\frac{3\ln(1+\frac{1}{4x})}{\frac{1}{x}}\] then use l'hopital

OpenStudy (anonymous):

would it be